# Mutual Information Density of Massive MIMO Systems over Rayleigh-Product Channels

### 05 Jan 2024 v4

@@ Appendix C Proof of Theorem 2
${\mathbb{A}}={\mathbb{A}}_{n}$ $={\mathbb{I}}_{N}-\frac{{\mathbb{X}}^{(n)}({\mathbb{X}}^{(n)})^{H}}{n},$ (83)
@@ C-A Step:1 Expectation over ${\mathbb{W}}$
$\frac{\partial\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}(u)}{\partial u}$ $=\mathbb{E}[(\jmath\mu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}-u\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ (88)
$+{\mathcal{O}}(\frac{u^{2}}{M}+\frac{u^{3}\mathcal{M}(u)}{M^{2}}+\frac{u^{4}\mathcal{M}(u)}{M^{3}}),$
@@
$\mu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ $=\sqrt{\frac{n}{M}}\log\det\left(\mathbb{I}_{N}+{\mathbb{H}}{\mathbb{H}}^{H}\right)-\frac{n}{\sqrt{Mn}}\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}),$ (91)
$\nu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ $=\frac{n}{Mn}\operatorname{Tr}(({\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})^{2})+\frac{2\sigma^{2}n}{Mn}\operatorname{Tr}({\mathbb{Q}}^{2}{\mathbb{H}}\frac{{\mathbb{X}}{\mathbb{X}}^{H}}{n}{\mathbb{H}}^{H}),$
@@
$+\frac{3\sigma^{2}n}{\sqrt{n^{3}M^{3}}}\operatorname{Tr}({\mathbb{Q}}^{2}{\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}\frac{{\mathbb{X}}{\mathbb{X}}^{H}}{n}{\mathbb{H}}^{H}),$
@@
where $\varepsilon_{w}(u)=\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}(u)\int_{0}^{u}e^{-x\alpha^{{\mathbb{Y}},{\mathbb{Z}}}_{n}}{\mathcal{O}}(\frac{x^{2}}{M^{2}})\mathrm{d}x]$ can be bounded by
@@
$=\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}(\frac{u\mathcal{M}(u)}{M^{2}}).$
@@
C-B Step 2: Evaluate $\mathbb{E}(\jmath\mu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}-u\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$
$\mathbb{E}(\jmath\mu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}-u\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ can be written as
$\mathbb{E}(\jmath\mu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}-u\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n})e^{\alpha^{{\mathbb{Y}},{\mathbb{Z}}}_{n}}=\mathbb{E}\jmath\mu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}-u\mathbb{E}\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=U_{1}+U_{2}.$ (96)
@@ C-B1 The evaluation of $U_{1}$
$U_{1}=(\jmath\sqrt{\frac{n}{M}}\int_{\sigma^{2}}^{\infty}\frac{N\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{z}-\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z)\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}\mathrm{d}z)-\frac{n}{\sqrt{Mn}}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=U_{1,1}+U_{1,2}.$ (97)
@@
In the following, we will evaluate $U_{1,1}$ and $U_{1,2}$, respectively. We first evaluate $\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z)\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ in $U_{1,1}$, which can be converted to the evaluation of $\mathbb{E}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ by the resolvent identity ${\mathbb{I}}_{N}=z{\mathbb{Q}}(z)+{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}$. By the integration by parts formula in (55), we can obtain
$\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=\mathbb{E}Y_{j,i}^{*}[{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}]_{j,i}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=\mathbb{E}[\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}-\frac{\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}}{M}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}$
$+\frac{\jmath u\sqrt{n}}{\sqrt{M^{3}}}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}-\frac{\jmath un}{\sqrt{M^{3}n}}\operatorname{Tr}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}+\frac{\jmath un}{\sqrt{M^{3}n}}\operatorname{Tr}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}$
@@
$-\sum_{i,j}\frac{u^{2}}{2M}\frac{\partial\nu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}[{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}]_{j,i}+\sum_{i,j}\frac{\jmath u^{3}}{3M}\frac{\partial\beta_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}[{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}]_{j,i}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}\overset{(a)}{=}\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}}{M}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$
$+\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}(1-\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}}{M})-\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}}{M}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$
$+\frac{\jmath u\sqrt{n}}{\sqrt{M^{3}}}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\mathbb{E}\varepsilon_{1}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\mathbb{E}\varepsilon_{2}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\mathbb{E}\varepsilon_{3}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\mathbb{E}\varepsilon_{4}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$
$+{\mathcal{O}}(\frac{\mathcal{P}_{1}(\frac{1}{z})}{zM}+\frac{u\mathcal{P}_{2}(\frac{1}{z})\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{3}{2}}}+\frac{u^{2}\mathcal{P}_{3}(\frac{1}{z})(1+\sqrt{\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}})}{M^{2}}+\frac{u^{3}\mathcal{P}_{4}(\frac{1}{z})(1+\sqrt{\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}})}{M^{3}})$ (98)
@@
=$\overset{(b)}{=}M\omega^{2}_{z}\overline{\omega}^{2}_{z}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\overline{\omega}^{2}_{z}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\frac{\jmath u\overline{\omega}_{z}\sqrt{n}}{\sqrt{M^{3}}}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$
$+\overline{\omega}_{z}\mathbb{E}\varepsilon_{1}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\overline{\omega}_{z}\mathbb{E}\varepsilon_{2}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\overline{\omega}_{z}\mathbb{E}\varepsilon_{3}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\overline{\omega}_{z}\mathbb{E}\varepsilon_{4}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$
@@
$+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{z}(\frac{1}{M}+\frac{u\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{3}{2}}}+\frac{u^{2}(1+\sqrt{\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}})}{M^{2}}+\frac{u^{3}(1+\sqrt{\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}})}{M^{3}}),$
@@
$\varepsilon_{1}=-\frac{\jmath un}{\sqrt{M^{3}n}}\operatorname{Tr}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}},$
$\varepsilon_{2}=\frac{\jmath un}{\sqrt{M^{3}n}}\operatorname{Tr}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}},$
@@
$\mathbb{E}ab\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ $=\mathbb{E}a\mathbb{E}b\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\mathbb{E}\underline{a}b\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ (100)
$=\mathbb{E}a\mathbb{E}b\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\mathbb{E}b\mathbb{E}a\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}-\mathbb{E}a\mathbb{E}b\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\mathbb{E}\underline{a}{\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}\underline{b}$
$=\mathbb{E}a\mathbb{E}b\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\mathbb{E}b\mathbb{E}a\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}-\mathbb{E}a\mathbb{E}b\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{Mz^{2}}),$
where $a=\frac{\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}}{M}$, $b=\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}$. The term ${\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{Mz^{2}})$ follows from (58) and (59) of Lemma 2 in Appendix B with
$|\mathbb{E}\underline{a}{\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}\underline{b}|\leq\mathrm{Var}^{\frac{1}{2}}(a)\mathrm{Var}^{\frac{1}{2}}(b)={\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{Mz^{2}}).$ (101)

### 09 Apr 2024 v5

@@ Appendix C Proof of Theorem 2
${\mathbb{A}}={\mathbb{A}}_{n}$ $={\mathbb{I}}_{M}-\frac{{\mathbb{X}}^{(n)}({\mathbb{X}}^{(n)})^{H}}{n},$ (83)
@@ C-A Step:1 Expectation over ${\mathbb{W}}$
$\frac{\partial\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}(u)}{\partial u}$ $=\mathbb{E}[(\jmath\mu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}-u\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}\Bigl{(}\frac{u^{2}}{M}+\frac{u^{3}\mathcal{M}(u)}{M^{2}}+\frac{u^{4}\mathcal{M}(u)}{M^{3}}\Bigr{)},$ (88)
@@
$\mu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ $=\sqrt{\frac{n}{M}}\log\det\left(\mathbb{I}_{N}+\frac{1}{\sigma^{2}}{\mathbb{H}}{\mathbb{H}}^{H}\right)-\frac{n\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})}{\sqrt{Mn}},$ (91)
$\nu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ $=\frac{n}{Mn}\operatorname{Tr}(({\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})^{2})+\frac{2\sigma^{2}n}{Mn}\operatorname{Tr}\Bigl{(}{\mathbb{Q}}^{2}{\mathbb{H}}\frac{{\mathbb{X}}{\mathbb{X}}^{H}}{n}{\mathbb{H}}^{H}\Bigr{)},$
@@
$+\frac{3\sigma^{2}n}{\sqrt{n^{3}M^{3}}}\operatorname{Tr}\Bigl{(}{\mathbb{Q}}^{2}{\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}\frac{{\mathbb{X}}{\mathbb{X}}^{H}}{n}{\mathbb{H}}^{H}\Bigr{)},$
@@
where $\varepsilon_{w}(u)=\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}\int_{0}^{u}e^{-x\alpha^{{\mathbb{Y}},{\mathbb{Z}}}_{n}}{\mathcal{O}}(\frac{x^{2}}{M^{2}})\mathrm{d}x]$ can be bounded by
@@
$=\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}\Bigl{(}\frac{u\mathcal{M}(u)}{M^{2}}\Bigr{)}.$
@@
C-B Step 2: Expectation Over ${\mathbb{Z}}$ and ${\mathbb{Y}}$: Evaluation of $\mathbb{E}[(\jmath\mu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}-u\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$
In this step, we will utilize Gaussian tool, i.e., integration by parts formula [19, Eq. (17)] and the variance control in Lemma 2 to compute the expectation over ${\mathbb{Z}}$ and ${\mathbb{Y}}$ to prove (87). To this end, we first decompose $\mathbb{E}[(\jmath\mu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}-u\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ into $U_{1}$ and $U_{2}$ as
$\mathbb{E}[(\jmath\mu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}-u\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n})e^{\alpha^{{\mathbb{Y}},{\mathbb{Z}}}_{n}}]=\mathbb{E}[\jmath\mu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]-u\mathbb{E}[\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=U_{1}+U_{2}.$ (96)
@@ C-B1 The evaluation of $U_{1}$
$U_{1}=(\jmath\sqrt{\frac{n}{M}}\int_{\sigma^{2}}^{\infty}\frac{N\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]}{z}-\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z))\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]\mathrm{d}z)-\frac{n}{\sqrt{Mn}}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=U_{1,1}+U_{1,2}.$ (97)
@@
In the following, we will evaluate $U_{1,1}$ and $U_{1,2}$, respectively. We first evaluate $\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z))\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ in $U_{1,1}$, which can be converted to the evaluation of $\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ by the resolvent identity ${\mathbb{I}}_{N}=z{\mathbb{Q}}(z)+{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}$. By the integration by parts formula in (55), we can obtain
$\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=\mathbb{E}[Y_{j,i}^{*}[{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}]_{j,i}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=\mathbb{E}\Bigl{[}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})-\frac{\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})}{M}\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})$
$+\frac{\jmath u\sqrt{n}}{\sqrt{M^{3}}}\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})-\frac{\jmath un}{\sqrt{M^{3}n}}\operatorname{Tr}({\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}})+\frac{\jmath un}{\sqrt{M^{3}n}}\operatorname{Tr}({\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}})$
@@
$-\sum_{i,j}\frac{u^{2}}{2M}\frac{\partial\nu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}[{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}]_{j,i}+\sum_{i,j}\frac{\jmath u^{3}}{3M}\frac{\partial\beta_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}[{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}]_{j,i}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}\Bigr{]}\overset{(a)}{=}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$
$+\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}](1-\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})]}{M})-\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$
$+\frac{\jmath u\sqrt{n}}{\sqrt{M^{3}}}\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\mathbb{E}[\varepsilon_{1}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\mathbb{E}[\varepsilon_{2}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\mathbb{E}[\varepsilon_{3}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\mathbb{E}[\varepsilon_{4}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$
$+{\mathcal{O}}\Biggl{(}\frac{\mathcal{P}_{1}(\frac{1}{z})}{zM}+\frac{u\mathcal{P}_{2}(\frac{1}{z})\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{M^{\frac{3}{2}}}+\frac{u^{2}\mathcal{P}_{3}(\frac{1}{z})(1+\sqrt{\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}})}{M^{2}}+\frac{u^{3}\mathcal{P}_{4}(\frac{1}{z})(1+\sqrt{\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}})}{M^{3}}\Biggr{)}$ (98)
@@
=$\overset{(b)}{=}M\omega^{2}_{z}\overline{\omega}^{2}_{z}\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\overline{\omega}^{2}_{z}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\frac{\jmath u\overline{\omega}_{z}\sqrt{n}}{\sqrt{M^{3}}}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$
$+\overline{\omega}_{z}\mathbb{E}[\varepsilon_{1}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\overline{\omega}_{z}\mathbb{E}[\varepsilon_{2}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\overline{\omega}_{z}\mathbb{E}[\varepsilon_{3}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\overline{\omega}_{z}\mathbb{E}[\varepsilon_{4}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$
@@
$+{\mathcal{O}}\Biggl{(}\frac{\mathcal{P}(\frac{1}{z})}{z}\Bigl{(}\frac{1}{M}+\frac{u\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{M^{\frac{3}{2}}}+\frac{u^{2}(1+\sqrt{\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}})}{M^{2}}+\frac{u^{3}(1+\sqrt{\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}})}{M^{3}}\Bigr{)}\Biggr{)},$
@@
$\varepsilon_{1}$ $=-\frac{\jmath un}{\sqrt{M^{3}n}}\operatorname{Tr}({\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}),$
$\varepsilon_{2}$ $=\frac{\jmath un}{\sqrt{M^{3}n}}\operatorname{Tr}({\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}),$
@@
$\mathbb{E}[ab\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ $=\mathbb{E}[a]\mathbb{E}[b\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\mathbb{E}[\underline{a}b\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ (100)
$=\mathbb{E}[a]\mathbb{E}[b\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\mathbb{E}[b]\mathbb{E}[a\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]-\mathbb{E}[a]\mathbb{E}[b]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\mathbb{E}[\underline{a}{\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}\underline{b}]$
$=\mathbb{E}[a]\mathbb{E}[b\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\mathbb{E}[b]\mathbb{E}[a\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]-\mathbb{E}[a]\mathbb{E}[b]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{Mz^{2}}),$
where $a=\frac{\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})}{M}$, $b=\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})$. The term ${\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{Mz^{2}})$ follows from (58) and (59) of Lemma 2 in Appendix B with
$|\mathbb{E}[\underline{a}{\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}\underline{b}]|\leq\mathrm{Var}^{\frac{1}{2}}(a)\mathrm{Var}^{\frac{1}{2}}(b)={\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{Mz^{2}}\Bigr{)}.$ (101)